# ECORE: Energy-Conscious Optimized Routing for Deep Learning Models at the Edge

### v3

@@ I Introduction
Our recent findings show that detection models exhibit markedly different energy, latency, and accuracy profiles when deployed diverse edge devices [5]. By leveraging this hardware–model diversity, systems can achieve substantial energy savings without compromising detection accuracy or latency. Energy efficiency is particularly critical on edge devices operating under strict power constraints, such as those relying on batteries or local energy harvesting methods like solar panels. Yet in safety-critical applications like surveillance and traffic monitoring, accuracy remains non-negotiable. This creates the central challenge addressed in this study: “how can we minimize energy usage without degrading detection quality?
@@
ECORE’s routers use lightweight estimation techniques during the pre-processing stage to infer the number of objects in each incoming image as an indicator of scene complexity. To the best of our knowledge, this is the first work to address the routing problem through the lens of scene complexity. Based on this estimate, the router selects and forwards the request to an optimal model–device pair. The main contributions of this work are as follows:
@@
    A first-of-its-kind system architecture for a framework called ECORE , providing optimized routing of image processing tasks across heterogeneous edge devices running diverse object detection models, coordinated by a central gateway responsible for routing decisions.
@@
    Three novel object count estimation approaches to support intelligent routing decisions in ECORE . The first is a light-weight edge detection–based approach, which utilizes the Canny technique to estimate the number of objects in an image. The second is an SSD-based front-end approach, which deploys a minimal object detection model at the gateway to make decisions. The third is low-overhead approach approach, which leverages the object detection outputs returned from the backbend in the previous processing round and uses this information to inform routing decisions for the subsequent request.
@@
    An extensive experimental evaluation of our proposed routing approaches on a real-world testbed, comparing them against an Oracle Router (which has perfect knowledge of the exact number of objects in each image) as well as several baseline techniques, including Random, Round Robin, Lowest Energy Consumption, Lowest Inference Time, and Highest mean Average Precision (mAP). The evaluation leverages multiple realistic datasets, such as the Common Objects in Context (COCO) validation set and a practical surveillance video stream, tested under diverse input scenarios.
The remainder of the paper is organized as follows:Section II discusses the motivation behind this work. Section III presents the system design, problem formulation, and greedy routing approach. Section IV describes the datasets, experimental setup, and results. Section V reviews related literature. Finally, Section VI concludes the paper and outlines directions for future work.
@@ II Motivation
To verify our hypothesis, we conducted a preliminary experiment using the COCO validation dataset, which contains 5,000 images. We categorized the images into two groups based on the number of objects in the image, as determined by the ground truth: one group with a single object per image and another with four or more. We then calculated the mAP and energy consumption per image during inference. As shown in Fig. 2(b)(b), SSD Lite and YOLOv8n exhibit similar accuracy for single-object images, while YOLOv8n significantly outperforms SSD Lite on images with four or more objects, achieving nearly twice better mAP. In terms of energy usage, SSD Lite consumes a consistent amount across both groups, which is approximately 50% lower than YOLOv8n. These results suggest that in less complex scenarios, two different object detection models can achieve similar accuracy, while significant energy savings is possible by using more efficient models such as SSD Lite without compromising accuracy.
@@ III System Design and Implementation
This section describes the system architecture of ECORE , designed for energy-efficient and performance-aware routing of object detection tasks in an edge environment. Fig. 2 illustrates the overall design. We assume a scenario where a pool of backend edge devices hosts a variety of model–device pairs. Each device runs a dedicated object detection model and exposes an inference API to process images or video frames captured via distributed cameras or embedded sensors.
Cameras or embedded sensors transmit image data or video frames to a central gateway, which performs lightweight preprocessing and analysis. As shown in Fig. 2, the gateway first applies one of the proposed object estimation techniques to infer the number of objects in each incoming image. This count is passed to the routing algorithm (Algorithm 1), which uses it to determine the most appropriate model–device pair.
The routing algorithm begins by identifying the corresponding object count group based on the estimated number of objects, using a predefined set of group rules defined by numeric ranges and labels. It then filters the profiling data to retain only the entries that match the identified group. From this subset, it calculates the maximum achievable mAP and derives a minimum acceptable mAP within a predefined tolerance margin. The data is then filtered again to retain only model–device pairs whose mAP falls within this target range. Among these, the pair with the lowest energy consumption is selected. The gateway forwards the request to the selected endpoint, which processes the image and returns the detected objects and their bounding boxes.
@@ III-A Problem Formulation and Greedy Approach
When running an object detection application in an edge environment, a key challenge lies in routing image requests to the most suitable model-device pair, taking into account trade-offs among accuracy, latency, and energy. Let:
@@
Given an incoming image request accompanied by an estimated object count group $G$, the routing objective selects the most energy-efficient model–device pair that satisfies application constraints for accuracy. This decision-making process can be formulated as a constrained optimization problem. First, the subset $\mathcal{M}_{G}\subseteq\mathcal{M}$ is identified, containing only the pairs belonging to group $G$. From this subset, the algorithm determines the maximum achievable accuracy.
@@
To preserve flexibility in model-device selection and allow for more energy-efficient decision, we introduce a dynamic tolerance margin, $\delta_{\text{mAP}}$. This controlled relaxation enables the system to trade off a small reduction in mAP for gains in energy efficiency. The margin defines a feasible set $\mathcal{F}\subseteq\mathcal{M}_{G}$, within which model-device pairs are considered acceptable for routing, and formally defined as:
@@ III-B Optimality of the Routing Algorithm
Let $\mathcal{M}_{G}$ be the subset of profiled model-device pairs matching the estimated object count group $G$, and let $\delta_{\text{mAP}}$ be a fixed accuracy tolerance. Then, the greedy selection of the model-device pair with the lowest energy consumption from the filtered feasible set
@@
where the feasible set $\mathcal{F}$ includes only model-device pairs in $\mathcal{M}_{G}$ whose accuracy exceeds the threshold $\text{mAP}_{\max}-\delta_{\text{mAP}}$.
@@ III-C Proposed object count estimation approaches
To avoid negating energy savings, the routing algorithm must remain lightweight, especially its object-counting component. We therefore design and evaluate three approaches for estimating the number of objects in incoming image requests and directing them to the most suitable edge node.”
  1. 1.
    Edge Detection (ED): This approach leverages edge detection which is a fundamental technique in image processing and computer vision used to identify areas of significant intensity or color change, which often correspond to object boundaries. The primary goal is to capture essential structural information. We employ the widely used Canny edge detection algorithm [6], to estimate the number of objects in each image. The resulting object count is then fed into the proposed routing algorithm (Algorithm 1), which selects the optimal model-device pair based on the expected energy–accuracy trade-off governed by $\delta_{\text{mAP}}$. This method is considered lightweight and computationally efficient, though it may yield coarse or approximate estimations.
  2. 2.
@@
  3. 3.
    Output-Based (OB): This approach leverages temporal continuity in image streams, particularly video, by assuming that consecutive frames contain a similar number of objects. Instead of estimating the object count for every image (frame), the system reuses the count obtained from the output of the previous image (frame), as detected by the last selected model–device pair. The same configuration is retained until a change in object count is observed. The routing algorithm begins with a default estimate (e.g., zero) and forwards the first image to an edge node. Once the detection results are returned, the router extracts the actual number of objects and uses it for subsequent routing decisions. If the object count remains stable, the system continues using the same model–device pair; if a change is detected, the routing decision is updated accordingly. This adaptive strategy avoids per-image (frame) estimation, reducing computational overhead and improving energy efficiency.
@@ IV-A2 Testbed
This section outlines the experimental testbed used to evaluate the proposed routing approach. The testbed configuration builds on exhaustive prior benchmarking results [5], which involved a comprehensive and resource-intensive evaluation of various combinations of edge devices and object detection models across three key performance metrics: energy consumption, inference time, and mAP. This benchmarking effort, which profiled a wide range of devices and models, provides an invaluable reference point for the research community studying edge-based object detection. Specifically, the devices evaluated included Raspberry Pi 3 (with and without TPU), Raspberry Pi 4 (with and without TPU), Raspberry Pi 5 (with and without TPU and with an AI HAT), and the Jetson Orin Nano, making a total of eight. The eight profiled object detection models spanned from lightweight variants—SSD v1, SSD Lite, EfficientDet-Lite0, Lite1, and Lite2—to the YOLOv8 family (Nano, Small, and Medium). Together, these metrics reflect the inherent trade-offs in deploying object detection models on edge hardware. Notably, the benchmarking revealed that no single model–device pair consistently outperforms others across all criteria, reinforcing the need for context-dependent model selection guided by specific performance objectives. Fig. 4 illustrates the accuracy–energy consumption trade-offs for all 64 model–device combinations.
@@
To construct our testbed, we selected only model–device pairs that lie on or near the Pareto front in at least one performance dimension. We repeated all the configurations in Fig. 4 for five object groups—‘0’, ‘1’, ‘2’, ‘3’, and ‘4 or more’ objects per image—an effort that required extensive experimentation and significant computational time. This ensures that all included devices offer meaningful trade-offs—i.e., each is strong in at least one metric—while excluding configurations that are consistently suboptimal across all criteria. Energy efficiency and latency do not vary across groups. Therefore, for energy efficiency, we selected Jetson Orin Nano running SSD v1 which demonstrated the lowest energy consumption among all configurations. In terms of inference latency, the Raspberry Pi 5 with TPU paired with SSD v1 achieved the shortest inference time. For detection accuracy (mAP), the optimal configuration varied across object count groups, underscoring the need for adaptive routing. Specifically, for Group 1 (images with no objects), the Raspberry Pi 5 with TPU running SSD v1 achieved the highest mAP. For Group 2 (images with one object), the Raspberry Pi 5 with TPU running SSD Lite performed best. In Group 3 (images with two objects), the Jetson Orin Nano with YOLOv8-small yielded the highest accuracy. Finally, for Groups 4 and 5 (images with three and four or more objects), the Raspberry Pi 5 equipped with an AI Hat running YOLOv8-small demonstrated superior performance. Table I provides the finalized list of device-model pairs, and Table II details the corresponding device specifications. This testbed supports a thorough assessment of the routing algorithm’s effectiveness in achieving optimal trade-offs between energy efficiency and performance across a variety of edge computing environments.
@@
Metrics Edge Device Object Detection Model Runtime
Energy Consumption Jetson Orin Nano SSD v1 TensorRT
@@
Device Name Processor Memory OS/SDK Bluetooth Wi-Fi
Raspberry Pi 4 CPU 4 GB Debian Bookworm Off Off
Pi 5 + Coral TPU TPU 4 GB Debian Bookworm Off Off
Pi 5 + Coral TPU TPU 4 GB Debian Bookworm Off Off
Pi 5 + AI Hat Hailo-8 4 GB Debian Bookworm Off Off
Jetson Orin Nano GPU 8 GB JetPack 5.1.3 Off Off
Jetson Orin Nano GPU 8 GB JetPack 5.1.3 Off Off
@@ IV-B Experimental Procedure
To the best of our knowledge, no existing work in the literature directly addresses the same problem with a comparable approach. Therefore, to rigorously evaluate the effectiveness of our proposed routing strategies, we systematically compared them against a set of baseline routing methods, including an idealized benchmark referred to as the Oracle Router, as well as several widely used strategies: Round Robin, Random, Lowest Energy, Lowest Inference Time, Highest mAP an Highest mAP per Group. The baseline and benchmark routing strategies are defined as follows:
@@
  • Random Router (Rnd): Assigns each request to a node randomly, independent of any system metrics or workload history.
  • Lowest Energy Router (LE): Routes each image to the model–device pair that consumes the least energy, prioritizing power efficiency over accuracy or latency.
@@
  • Highest mAP Router (HM): Always chooses the model–device pair with the highest mAP, independent of the object count in the image.
  • Highest mAP per Group Router (HMG): Within each object-count group, the model–device pair achieving the highest mAP is selected, irrespective of energy consumption or latency constraints.
@@
To ensure consistent load generation, we used Locust  [^1] https://locust.io/ to send image requests back-to-back in a piggybacked fashion where each new request is triggered only after the previous image has been processed and a response is received, simulating continuous workloads. Performance was evaluated using the following primary metrics:
@@ IV-C1 Full COCO Dataset Results:
Insight #1. Considering the upper bound on achievable mAP (as given by HMG) and the minimum energy consumption (as achieved by LE), the proposed ED approach offers the most efficient trade-off on static-image datasets such as COCO. It attains accuracy close to the oracle-level performance—which has full knowledge of object counts—while maintaining competitive energy efficiency and incurring only minimal latency overhead.
[…]
@@ IV-D Discussions
ECORE demonstrates promising results by reducing energy consumption and latency while maintaining strong accuracy. While the Orc router represents an idealized baseline, benefiting from prior knowledge of object counts, achieving comparably strong results in practice is not infeasible. Our proposed routing approaches—ED and OB—achieve performance close to Oracle, with marginal differences. However, the effectiveness of each method is dataset-dependent: OB performs optimally in sequential and structured scenarios, such as the sorted balanced and video datasets, while ED is better suited for static image inputs without any continuity.
The current design assumes static profiling, where performance metrics remain stable at runtime, which may not reflect real-world dynamics such as temperature, background load, or battery state. Additionally, our heuristic operates at single-request granularity, limiting its applicability in batch or load-balancing contexts. It is important to note that the proposed approaches do not support multi-objective optimization; balancing trade-offs between energy and latency would require a Pareto-optimal or weighted strategy, where a greedy algorithm may no longer suffice.

### v4

@@ I Introduction
Our recent findings show that detection models exhibit markedly different energy, latency, and accuracy profiles when deployed on diverse edge devices [2]. By leveraging this hardware–model diversity, systems can achieve substantial energy savings without compromising detection accuracy or latency. Energy efficiency is particularly critical on edge devices operating under strict power constraints, such as those relying on batteries or local energy harvesting methods like solar panels. Yet in safety-critical applications like surveillance and traffic monitoring, accuracy remains non-negotiable. This creates the central challenge addressed in this study: “how can we minimize energy usage without degrading detection quality?
@@
ECORE’s routers use lightweight estimation techniques during the pre-processing stage to infer the number of objects in each incoming image as an indicator of scene complexity. To the best of our knowledge, this is the first work to address the routing problem through the lens of scene complexity. Based on this estimate, the router selects and forwards the request to an optimal device-model pair. The main contributions of this work are as follows:
@@
    We propose a novel system architecture, ECORE , that optimizes the routing of image processing tasks across heterogeneous edge devices running diverse object detection models, coordinated by a central gateway.
@@
    We introduce three object count estimation techniques to enable intelligent routing in ECORE : (i) a lightweight edge-detection–based method using Canny, (ii) an SSD-based lightweight detection model deployed at the gateway, and (iii) a low-overhead approach that exploits object detection results from previous inference rounds.
@@
    We conduct an extensive experimental evaluation on a real-world testbed, comparing our approaches against an Oracle Router and multiple baselines using realistic datasets including COCO and a surveillance video stream under diverse scenarios.
@@ II Motivation
To verify our hypothesis, we conducted a preliminary experiment using the COCO validation dataset, which contains 5,000 images. We categorized the images into two groups based on the number of objects in the image, as determined by the ground truth: one group with a single object per image and another with four or more. We then calculated the mAP and energy consumption per image during inference. As shown in Fig. 2(b)(b), SSD Lite and YOLOv8 small exhibit similar accuracy for single-object images, while YOLOv8 small significantly outperforms SSD Lite on images with four or more objects, achieving nearly twice better mAP. In terms of energy usage, SSD Lite consumes a consistent amount across both groups, which is approximately 50% lower than YOLOv8 small. These results suggest that in less complex scenarios, two different object detection models can achieve similar accuracy, while significant energy savings is possible by using more efficient models such as SSD Lite without compromising accuracy.
@@ III System Design and Implementation
This section describes the system architecture of ECORE , designed for energy-efficient and performance-aware routing of object detection tasks in an edge environment. Fig. 2 illustrates the overall design. We assume a scenario where a pool of backend edge devices hosts a variety of device-model pairs. Each device runs a dedicated object detection model and exposes an inference API to process images or video frames captured via distributed cameras or embedded sensors. Cameras or embedded sensors transmit image data or video frames to a central gateway, which performs lightweight preprocessing and request coordination. This gateway-centric design reflects common smart-surveillance deployments (e.g., CCTV), decoupling cameras from device heterogeneity and centralizing monitoring and decision-making. The same workflow also applies to other pervasive edge settings where sensors generate inference requests and multiple heterogeneous edge nodes are available (e.g., smart retail, traffic monitoring, industrial inspection, and campus IoT). As shown in Fig. 2, the gateway applies one of the proposed object estimation techniques to infer the number of objects in each incoming image. This object-estimation component is modular, and alternative estimators can be substituted without changing the rest of the routing pipeline. The estimated count is passed to the routing algorithm (Algorithm 1), which uses it to select the most appropriate device-model pair.
The routing algorithm begins by identifying the corresponding object count group based on the estimated number of objects, using predefined group rules defined by numeric ranges and labels. It then filters the profiling data to retain only the entries that match the identified group. From this subset, it calculates the maximum achievable mAP and derives a minimum acceptable mAP within a predefined tolerance margin. The data is then filtered again to retain only device-model pairs whose mAP falls within this target range. Among these, the pair with the lowest energy consumption is selected. The gateway forwards the request to the selected endpoint, which processes the image and returns the detected objects and their bounding boxes.
@@ III-A Problem Formulation and Greedy Approach
When running an object detection application in an edge environment, a key challenge lies in routing image requests to the most suitable device-model pair, taking into account trade-offs among accuracy, latency, and energy. Let:
@@
Given an incoming image request accompanied by an estimated object count group $G$, the routing objective selects the most energy-efficient device-model pair that satisfies application constraints for accuracy. This decision-making process can be formulated as a constrained optimization problem. First, the subset $\mathcal{M}_{G}\subseteq\mathcal{M}$ is identified, containing only the pairs belonging to group $G$. From this subset, the algorithm determines the maximum achievable accuracy.
@@
To preserve flexibility in device-model selection and allow for more energy-efficient decision, we introduce a dynamic tolerance margin, $\delta_{\text{mAP}}$. This controlled relaxation enables the system to trade off a small reduction in mAP for gains in energy efficiency. The margin defines a feasible set $\mathcal{F}\subseteq\mathcal{M}_{G}$, within which device-model pairs are considered acceptable for routing, and formally defined as:
@@ III-B Optimality of the Routing Algorithm
Let $\mathcal{M}_{G}$ be the subset of profiled device-model pairs matching the estimated object count group $G$, and let $\delta_{\text{mAP}}$ be a fixed accuracy tolerance. Then, the greedy selection of the device-model pair with the lowest energy consumption from the filtered feasible set
@@
where the feasible set $\mathcal{F}$ includes only device-model pairs in $\mathcal{M}_{G}$ whose accuracy exceeds the threshold $\text{mAP}_{\max}-\delta_{\text{mAP}}$.
@@ III-C Proposed object count estimation approaches
To avoid negating energy savings, the routing algorithm must remain lightweight, especially its object-counting component. We therefore design and evaluate three approaches for estimating the number of objects in incoming image requests and directing them to the most suitable edge node.
–Edge Detection (ED): This approach leverages edge detection which is a fundamental technique in image processing and computer vision used to identify areas of significant intensity or color change, which often correspond to object boundaries. The primary goal is to capture essential structural information. We employ the widely used Canny edge detection algorithm [3], to estimate the number of objects in each image. The resulting object count is then fed into the proposed routing algorithm (Algorithm 1), which selects the optimal device-model pair based on the expected energy–accuracy trade-off governed by $\delta_{\text{mAP}}$. This method is considered lightweight and computationally efficient, though it may yield coarse or approximate estimations.
@@
–Output-Based (OB): This approach leverages temporal continuity in image streams, particularly video, by assuming that consecutive frames contain a similar number of objects. Instead of estimating the object count for every image (frame), the system reuses the count obtained from the output of the previous image (frame), as detected by the last selected device-model pair. The same configuration is retained until a change in object count is observed. The routing algorithm begins with a default estimate (e.g., zero) and forwards the first image to an edge node. Once the detection results are returned, the router extracts the actual number of objects and uses it for subsequent routing decisions. If the object count remains stable, the system continues using the same device-model pair; if a change is detected, the routing decision is updated accordingly. This adaptive strategy avoids per-image (frame) estimation, reducing computational overhead and improving energy efficiency.
@@ IV-A2 Testbed
This section outlines the experimental testbed used to evaluate the proposed routing approach. The testbed configuration builds on exhaustive prior benchmarking results [1], which involved a comprehensive and resource-intensive evaluation of various combinations of edge devices and object detection models across three key performance metrics: energy consumption, inference time, and mAP. This benchmarking effort, which profiled a wide range of devices and models, provides an invaluable reference point for the research community studying edge-based object detection. Specifically, the devices evaluated included Raspberry Pi 3 (with and without TPU), Raspberry Pi 4 (with and without TPU), Raspberry Pi 5 (with and without TPU and with an AI HAT), and the Jetson Orin Nano, making a total of eight. The eight profiled object detection models spanned from lightweight variants—SSD v1, SSD Lite, EfficientDet-Lite0, Lite1, and Lite2—to the YOLOv8 family (Nano, Small, and Medium). Together, these metrics reflect the inherent trade-offs in deploying object detection models on edge hardware. Notably, the benchmarking revealed that no single device-model pair consistently outperforms others across all criteria, reinforcing the need for context-dependent model selection guided by specific performance objectives. Fig. 4 illustrates the accuracy–energy consumption trade-offs for all 64 device-model combinations.
@@
To construct our testbed, we selected only device-model pairs that lie on or near the Pareto front in at least one performance dimension. We repeated all the configurations in Fig. 4 for five object groups—‘0’, ‘1’, ‘2’, ‘3’, and ‘4 or more’ objects per image—an effort that required extensive experimentation and significant computational time. This ensures that all included devices offer meaningful trade-offs—i.e., each is strong in at least one metric—while excluding configurations that are consistently suboptimal across all criteria. Energy efficiency and latency do not vary across groups. Therefore, for energy efficiency, we selected Jetson Orin Nano running SSD v1 which demonstrated the lowest energy consumption among all configurations. In terms of inference latency, the Raspberry Pi 5 with TPU paired with SSD v1 achieved the shortest inference time. For detection accuracy (mAP), the optimal configuration varied across object count groups, underscoring the need for adaptive routing. Specifically, for Group 1 (images with no objects), the Raspberry Pi 5 with TPU running SSD v1 achieved the highest mAP. For Group 2 (images with one object), the Raspberry Pi 5 with TPU running SSD Lite performed best. In Group 3 (images with two objects), the Jetson Orin Nano with YOLOv8-small yielded the highest accuracy. Finally, for Groups 4 and 5 (images with three and four or more objects), the Raspberry Pi 5 equipped with an AI Hat running YOLOv8-small demonstrated superior performance. Table I provides the finalized list of device-model pairs, and Table II details the corresponding device specifications. This testbed supports a thorough assessment of the routing algorithm’s effectiveness in achieving optimal trade-offs between energy efficiency and performance across a variety of edge computing environments.
@@
Metrics Edge Device Object Detection Runtime
Energy Cons. Jetson Orin Nano SSD v1 TensorRT
@@
Device Name Unit Mem. OS/SDK Bluetooth Wi-Fi
Pi 4 CPU 4 GB Debian Bookworm Off Off
Pi 5 + TPU TPU 4 GB Debian Bookworm Off Off
Pi 5 + TPU TPU 4 GB Debian Bookworm Off Off
Pi 5 + AI Hat NPU 4 GB Debian Bookworm Off Off
Orin Nano GPU 8 GB JetPack 5.1.3 Off Off
Orin Nano GPU 8 GB JetPack 5.1.3 Off Off
@@ IV-B Experimental Procedure
To the best of our knowledge, no existing work in the literature directly addresses the same problem with a comparable approach. Therefore, to rigorously evaluate the effectiveness of our proposed routing strategies, we systematically compared them against a set of baseline routing methods, as well as several widely used strategies:
@@
–Random Router (Rnd): Assigns requests to a node randomly, independent of any system metrics or workload history.
–Lowest Energy Router (LE): Routes each image to the device-model pair that consumes the least energy, prioritizing power efficiency over accuracy or latency.
@@
–Highest mAP Router (HM): Always chooses the device-model pair with the highest mAP, independent of the object count in the image.
–Highest mAP per Group Router (HMG): Within each object-count group, the device-model pair achieving the highest mAP is selected, irrespective of energy consumption.
@@
To ensure consistent load generation, we used Locust [^1] https://locust.io/ to send image requests back-to-back in a piggybacked fashion where each new request is triggered only after the previous image has been processed and a response is received, simulating continuous workloads. Performance is evaluated using energy consumption as the primary metric, with accuracy as a constraint, while latency and gateway overhead are reported as supporting quality-of-service (QoS) indicators:
@@ IV-C1 Full COCO Dataset Results:
Insight #1. Considering upper bound on achievable mAP (as given by HMG) and the minimum energy consumption (as achieved by LE), the proposed ED approach offers the most efficient trade-off on static-image datasets such as COCO. It attains accuracy close to the Orc performance-which has full knowledge of object counts-while maintaining competitive energy efficiency and incurring only minimal latency overhead.
@@ IV-D Discussions
ECORE demonstrates promising results by reducing energy consumption and latency while maintaining strong accuracy. While the Orc router represents an idealized baseline, benefiting from prior knowledge of object counts, achieving comparably strong results in practice is not infeasible. Our proposed routing approaches ED and OB achieve performance close to Oracle, with marginal differences. However, the effectiveness of each method is dataset-dependent: OB performs optimally in sequential and structured scenarios, such as the sorted balanced and video datasets, while ED is better suited for static image inputs without any continuity.
The current design assumes static profiling, where performance metrics remain stable at runtime, which may not reflect real-world dynamics such as temperature, background load, or battery state. To handle such dynamics, profiling can be updated online by periodically collecting recent measurements and refreshing the profiles, allowing ECORE to adapt to thermal throttling and background load. Additionally, our heuristic operates at single-request granularity, limiting its applicability in batch or load-balancing contexts. It is important to note that the proposed approaches do not support multi-objective optimization; balancing trade-offs between energy and latency would require a Pareto-optimal or weighted strategy, where a greedy algorithm may no longer suffice.